# Segment Anything Model for Medical Images?

### 12 Dec 2023 v5

@@ 3.3 Manual Prompt Mode
As shown in Fig. 5, our prompt mode contains five strategies, including one positive point ($S_{2}$), five positive points ($S_{3}$), five positive points with five negative points ($S_{4}$), one box ($S_{6}$), and one box with one positive point ($S_{6}$). We further established a unified rule for point selection to ensure randomness, repeatability, and accuracy. For the positive point selection, a) we first calculated the center of mass of the ground truth (GT) mask (red point in Fig. 5). b) If the center of mass was inside the GT mask, we took the center as the first positive point. c) Then, we directly flattened the GT mask to a one-dimensional vector and obtained the other positive points by adopting the uniform sampling method (green points in Fig. 5). d) If the center of mass was outside the GT mask, all the required positive points would be obtained by performing step c. For the negative point selection, we aimed to avoid selecting points that were too distant from the target region. Specifically, we first enlarged the bounding box of the GT by two times. The negative points were generated by even sampling in the non-GT region (yellow points in Fig. 5). Last, for the box selection, we directly adopted the bounding box of the GT mask without any additional operations. The above strategy can ensure the repeatability of the experiments. Besides, we tend to test the theoretical optimum performance of SAM via the selection of the center of mass and tight box. Since they may include the most representative features of the target. It is noted that SAM allows multiple prompts to be input into the network once. Thus, for a fair comparison, we test the one-round interaction performance of SAM under the above five prompt strategies ($S_{2}$-$S_{6}$).
@@ 4.2 Evaluation Metrics
    DICE Coefficient (DICE, %) (Bilic et al., 2023): A similarity measure to evaluate the overlap between the prediction and GT. Ranging from [0, 1], a higher value indicated a better performance of the model.
@@
    Jaccard Similarity Coefficient (JAC, %) (Crum et al., 2006): Also called IOU, is used to measure the similarity between two masks. It is similar to DICE, but with a different calculation method. Specifically, for the predicted and GT masks, A and B, JAC calculates the intersection ($|A\cap B|$) over union ($|A\cup B|$). JAC ranges from 0 to 1, and higher values indicate better performance.
@@
    Hausdorff Distance (HD, pixel) (Zhou et al., 2019a): A measure that evaluates the degree of similarity between two sets of points, which can reflect the distance between each point in the prediction to the points in the GT. It is more sensitive to the boundary than DICE.
@@ 4.7 Analysis on the Number of Points in Everything Mode
Adenocarcinoma-Histopathology $42.3_{44.1}$ $70.4_{37.4}$ $76.3_{32.5}$ $77.8_{31.1}$ $78.7_{30.4}$ $79.1_{30.2}$
Adenocarcinoma-Histopathology $55.6_{42.4}$ $70.9_{34.7}$ $74.2_{31.5}$ $76.3_{30.0}$ $77.1_{29.1}$ $77.6_{28.6}$
Mitochondria-Electron Microscopy $31.8_{39.3}$ $71.1_{31.4}$ $81.1_{20.1}$ $81.5_{19.1}$ $81.7_{18.6}$ $81.7_{18.3}$
Neural Structures-Microscopy $21.3_{38.0}$ $45.7_{44.1}$ $61.5_{40.1}$ $64.4_{38.4}$ $65.4_{37.8}$ $65.8_{37.5}$
@@ 4.8 Analysis of Factors Correlating to Segmentation Results
We conducted an analysis of the partial correlation between the five attributes of the target objects mentioned above and the DICE score using Spearman’s rank partial correlation coefficient (Williams, 2021) under different testing strategies. The statistical results are displayed in Table 6, while Fig. 13 illustrates the scatterplot for the $S_{5}$ strategies. Across most testing strategies, we observed that the DICE score exhibited a moderate correlation ($0.70.4\leq\rho<$ 0.7) with the FO and intensity difference, a weak correlation ($0.40.2\leq\rho<$ 0.4) with the size, and no correlation with the modality and aspect ratio. Thus, SAM can consistently segment medical targets with different modality and aspect ratios. The performance of SAM under box prompt might be affected by the size of anatomical structures. Furthermore, the performance of SAM tends to be poor under all testing strategies when dealing with objects characterized by complex boundaries or low contrast. To confirm these findings, we divided DICE calculated under $S_{5B}$ into ten levels on average (e.g., level-1 means DICE (%) belongs to (0,10]) and visualized the FO boxplot for different DICE levels in Fig. 14. The figure illustrates that as the DICE level increases, the FO distribution of the structures gradually shifts to a range with smaller values. Additionally, in Fig. 15, we presented visualizations of anatomical structures with various FO ranges. These visualizations reveal that the DICE score of anatomical structures tends to decrease as the FO increases. Further implies that shape and boundary complexity may have an impact on SAM’s segmentation performance.
@@ 4.11 Comparison between SAM and Interactive Methods
In the previous sections, we input all the prompts to the SAM’s prompt encoder once for a fair comparison of its one-round performance. To mimic the real-life interactive segmentation procedures, we performed multi-round SAM. The point selection strategy is similar to the common interactive methods. Specifically, SAM first clicks on the center of the target, then the remaining clicks are based on false negative (FN) and false positive (FP) regions. We then compared SAM with two different strong interaction segmentation approaches, i.e., FocalClick (Chen et al., 2022) and SimpleClick (Liu et al., 2022a). Both of them are pretrained on the same number of images as SAM was.

### 17 Jan 2024 v7

@@ 3.3 Manual Prompt Mode
As shown in Fig. 5, our prompt mode contains five strategies, including one positive point ($S_{2}$), five positive points ($S_{3}$), five positive points with five negative points ($S_{4}$), one box ($S_{5}$), and one box with one positive point ($S_{6}$). We further established a unified rule for point selection to ensure randomness, repeatability, and accuracy. For the positive point selection, a) we first calculated the center of mass of the ground truth (GT) mask (red point in Fig. 5). b) If the center of mass was inside the GT mask, we took the center as the first positive point. c) Then, we directly flattened the GT mask to a one-dimensional vector and obtained the other positive points by adopting the uniform sampling method (green points in Fig. 5). d) If the center of mass was outside the GT mask, all the required positive points would be obtained by performing step c. For the negative point selection, we aimed to avoid selecting points that were too distant from the target region. Specifically, we first enlarged the bounding box of the GT by two times. The negative points were generated by even sampling in the non-GT region (yellow points in Fig. 5). Last, for the box selection, we directly adopted the bounding box of the GT mask without any additional operations. The above strategy can ensure the repeatability of the experiments. Besides, we tend to test the theoretical optimum performance of SAM via the selection of the center of mass and tight box. Since they may include the most representative features of the target. It is noted that SAM allows multiple prompts to be input into the network once. Thus, for a fair comparison, we test the one-round interaction performance of SAM under the above five prompt strategies ($S_{2}$-$S_{6}$).
@@ 4.2 Evaluation Metrics
    DICE Coefficient (DICE, %) (Dice, 1945; Bilic et al., 2023): A similarity measure to evaluate the overlap between the prediction and GT. Ranging from [0, 1], a higher value indicated a better performance of the model.
@@
    Jaccard Similarity Coefficient (JAC, %) (Jaccard, 1908; Qian et al., 2022): Also called IOU, is used to measure the similarity between two masks. It is similar to DICE, but with a different calculation method. Specifically, for the predicted and GT masks, A and B, JAC calculates the intersection ($|A\cap B|$) over union ($|A\cup B|$). JAC ranges from 0 to 1, and higher values indicate better performance.
@@
    Hausdorff Distance (HD, pixel) (Crum et al., 2006; Karimi and Salcudean, 2020): A measure that evaluates the degree of similarity between two sets of points, which can reflect the distance between each point in the prediction to the points in the GT. It is more sensitive to the boundary than DICE.
@@ 4.7 Analysis on the Number of Points in Everything Mode
Adenocarcinoma-Histopathology (Sirinukunwattana et al., 2017) $42.3_{44.1}$ $70.4_{37.4}$ $76.3_{32.5}$ $77.8_{31.1}$ $78.7_{30.4}$ $79.1_{30.2}$
Adenocarcinoma-Histopathology (Graham et al., 2019) $55.6_{42.4}$ $70.9_{34.7}$ $74.2_{31.5}$ $76.3_{30.0}$ $77.1_{29.1}$ $77.6_{28.6}$
Mitochondria-Electron Microscopy (Lucchi et al., 2013) $31.8_{39.3}$ $71.1_{31.4}$ $81.1_{20.1}$ $81.5_{19.1}$ $81.7_{18.6}$ $81.7_{18.3}$
Neural Structures-Microscopy (Cardona et al., 2010) $21.3_{38.0}$ $45.7_{44.1}$ $61.5_{40.1}$ $64.4_{38.4}$ $65.4_{37.8}$ $65.8_{37.5}$
@@ 4.8 Analysis of Factors Correlating to Segmentation Results
We conducted an analysis of the partial correlation between the five attributes of the target objects mentioned above and the DICE score using Spearman’s rank partial correlation coefficient under different testing strategies. The statistical results are displayed in Table 6, while Fig. 13 illustrates the scatterplot for the $S_{5}$ strategies. Across most testing strategies, we observed that the DICE score exhibited a moderate correlation ($0.70.4\leq\rho<$ 0.7) with the FO and intensity difference, a weak correlation ($0.40.2\leq\rho<$ 0.4) with the size, and no correlation with the modality and aspect ratio. Thus, SAM can consistently segment medical targets with different modality and aspect ratios. The performance of SAM under box prompt might be affected by the size of anatomical structures. Furthermore, the performance of SAM tends to be poor under all testing strategies when dealing with objects characterized by complex boundaries or low contrast. To confirm these findings, we divided DICE calculated under $S_{5B}$ into ten levels on average (e.g., level-1 means DICE (%) belongs to (0,10]) and visualized the FO boxplot for different DICE levels in Fig. 14. The figure illustrates that as the DICE level increases, the FO distribution of the structures gradually shifts to a range with smaller values. Additionally, in Fig. 15, we presented visualizations of anatomical structures with various FO ranges. These visualizations reveal that the DICE score of anatomical structures tends to decrease as the FO increases. It further implies that shape and boundary complexity may have an impact on SAM’s segmentation performance.
@@ 4.11 Comparison between SAM and Interactive Methods
In the previous sections, we input all the prompts to the SAM’s prompt encoder once for a fair comparison of its one-round performance. To mimic the real-life interactive segmentation procedures, we performed multi-round SAM. The point selection strategy is similar to the common interactive methods. Specifically, SAM first clicks on the center of the target, then the remaining clicks are based on false negative (FN) and false positive (FP) regions. We then compared SAM with two different strong interaction segmentation approaches, i.e., FocalClick (Chen et al., 2022) and SimpleClick (Liu et al., 2023a). Both of them are pretrained on the same number of images as SAM was.
@@ 7 Acknowledgement
This work was supported by the grant from National Natural Science Foundation of China (Nos. 62171290, 62101343), Shenzhen-Hong Kong Joint Research Program (No. SGDX20201103095613036), and Shenzhen Science and Technology Innovations Committee (No. 20200812143441001).